Jefferson Lab HALL A Logbook for June 08 - All EXCEPT Automatic Entries
Main INDEX
Thursday
241948 06/19/08 20:37 folts BB electron detector
Tuesday
241946 06/17/08 10:18 Jack Detector gas systems shut down
Sunday
Saturday
Friday
241938 06/13/08 09:26 gilman carbon doors opened
241937 06/13/08 08:36 GILMAN CARBON DOORS
Thursday
241935 06/12/08 16:38 Kalyan Bigbite cosmic check
Wednesday
241871 06/11/08 13:35 folts target chamber vacuum
Tuesday
241870 06/10/08 18:51 xiaohui Correction for last entry
241869 06/10/08 18:45 xiaohui FPP carbon door
241859 06/10/08 17:03 brads FPP 3" doors looked OK on June 3
241851 06/10/08 15:39 gilman opened fpp doors
241850 06/10/08 13:59 Techs checklist
241849 06/10/08 09:41 folts Hall Access
Monday
241848 06/09/08 23:21 adaq Radiation and Ozone snapshot
241847 06/09/08 23:21 adaq Hall A Tools snapshot
241846 06/09/08 23:21 adaq BigBox Magnet snapshot
241845 06/09/08 23:21 adaq Beam Position and Current snapshot
241844 06/09/08 23:21 adaq Beam Energy in A and C snapshot
241842 06/09/08 22:25 V. Sulkosky Swing Shift Summary
241841 06/09/08 22:20 V. Sulkosky End of E08-007
241840 06/09/08 22:07 V. Sulkosky BigBite Magnet to Zero.
241835 06/09/08 21:54 Decowski Target moved to holey carbon
241824 06/09/08 21:25 xqian 700A 0.6uA 6cm LH2 BB DC HV
241819 06/09/08 21:20 xqian 700 A 1.1uA 6cm LH2 BB DC HV
241816 06/09/08 21:16 xqian 700 A 4.3uA 6cm LH2 BB DC HV
241814 06/09/08 21:13 xqian 700 A 3uA 6cm LH2 BB DC HV
241810 06/09/08 21:06 xqian 700 A 2uA BB DC HV
241807 06/09/08 20:56 xqian 0A BB magnet BBDC HV
241804 06/09/08 20:50 xqian 100 A BB magnet BB DC HV 2uA
241802 06/09/08 20:48 xqian 300A again
241801 06/09/08 20:47 xqian 2uA 300A on BB magnet BB DC HV
241795 06/09/08 20:34 xqian 500 A 2uA on LH2 BB DC HV
241790 06/09/08 20:24 adaq Radiation and Ozone snapshot
241789 06/09/08 20:24 adaq Hall A Tools snapshot
241788 06/09/08 20:24 adaq BigBox Magnet snapshot
241787 06/09/08 20:24 adaq Beam Position and Current snapshot
241786 06/09/08 20:24 adaq Beam Energy in A and C snapshot
241781 06/09/08 20:16 xqian 700 A Chamber dark current
241777 06/09/08 20:06 xqian HV screen shot for shower
241776 06/09/08 20:05 adaq Target Screens snapshot
241772 06/09/08 19:12 Decowski Target screens
241771 06/09/08 19:09 Decowski Target charts
241770 06/09/08 19:05 V. Sulkosky Hall A Tools and BCM
241765 06/09/08 17:46 Decowski Anomaly gone in run #2086
241762 06/09/08 17:35 V. Sulkosky LHRS HV GUI
241761 06/09/08 17:07 Decowski Anomaly in FPP Chamber 1 Wires
241758 06/09/08 16:25 V. Sulkosky Camera motion for LHRS Angle Stuck
241757 06/09/08 16:20 V. Sulkosky Spot Check 2085
241753 06/09/08 15:57 ellie Day Shift Summary
241752 06/09/08 15:32 R. Michaels CODA restart
241749 06/09/08 15:11 R. Michaels move to 6cm LH2
241747 06/09/08 15:04 xiaohui Carbon Door
241744 06/09/08 14:53 xiaohui HV map
241743 06/09/08 14:51 xiaohui Acceptance check K3 D+2%
241742 06/09/08 14:51 R. Michaels spot++ for run 2081
241741 06/09/08 14:50 xiaohui BB shower hit map
241739 06/09/08 14:49 R. Michaels move to Al dummy
241738 06/09/08 14:46 R. Michaels target screens, part II
241737 06/09/08 14:45 R. Michaels target screens, part I
241735 06/09/08 14:41 ellie L HRS Angle Screen Capture
241733 06/09/08 14:33 ellie Hall A Tools Screen Capture
241731 06/09/08 14:24 R. Michaels All LG on. Put in 6 cm LH2
241729 06/09/08 14:14 R. Michaels changes to kine 3
241728 06/09/08 13:23 ellie BB Magnet Screen Capture
241727 06/09/08 13:21 ellie BB HV Screen Capture
241726 06/09/08 13:19 adaq Spot++ of run 2079
241723 06/09/08 13:14 ellie BPM Screen
241722 06/09/08 13:09 adaq Hall A Tools Screen
241719 06/09/08 12:24 adaq Radiation and Ozone snapshot
241718 06/09/08 12:24 adaq Hall A Tools snapshot
241717 06/09/08 12:24 adaq BigBox Magnet snapshot
241716 06/09/08 12:23 adaq Beam Position and Current snapshot
241715 06/09/08 12:23 adaq Beam Energy in A and C snapshot
241714 06/09/08 12:05 adaq Target Screens snapshot
241707 06/09/08 09:58 brads Cerenkov HVs turned back on again
241706 06/09/08 09:55 xiaohui Hot wire in U1
241705 06/09/08 09:45 brad No HV data in Start Runs from 2058--2068
241704 06/09/08 09:39 ellie Current data matches reference (2)
241703 06/09/08 09:36 ellie Current run data matches reference
241700 06/09/08 09:10 ellie Screens of Abnormality in reference data
241699 06/09/08 09:07 adaq Spike in Reference Data
241693 06/09/08 07:55 xqian Owl shift summary
241692 06/09/08 07:37 xqian shift time counting
241689 06/09/08 06:58 xqian BB HV screen shot
241688 06/09/08 06:57 xqian run 2070 most of BB HV off
241687 06/09/08 06:55 xqian RUN 2069 K1 delta=2 BB HV all on
241682 06/09/08 06:12 Higinbotham Carbon Raster Off Run 2067
241669 06/09/08 04:23 adaq Radiation and Ozone snapshot
241668 06/09/08 04:23 adaq Hall A Tools snapshot
241667 06/09/08 04:23 adaq BigBox Magnet snapshot
241666 06/09/08 04:23 adaq Beam Position and Current snapshot
241665 06/09/08 04:23 adaq Beam Energy in A and C snapshot
241664 06/09/08 04:05 adaq Target Screens snapshot
241657 06/09/08 02:12 xqian a hot wire in u1
241656 06/09/08 02:11 xqian a shift in ch1 u hit
Sunday
241648 06/08/08 23:58 Guy Ron Swing Shift Summary
241645 06/08/08 23:31 jpchen Reduced loop 3 fan setting from 25% to 6%.
241634 06/08/08 20:23 adaq Radiation and Ozone snapshot
241633 06/08/08 20:23 adaq Hall A Tools snapshot
241632 06/08/08 20:23 adaq BigBox Magnet snapshot
241631 06/08/08 20:23 adaq Beam Position and Current snapshot
241630 06/08/08 20:23 adaq Beam Energy in A and C snapshot
241629 06/08/08 20:20 xiaohui Acceptance check
241628 06/08/08 20:19 xiaohui BB shower hit map
241626 06/08/08 20:10 Guy Ron BB HV Map
241624 06/08/08 20:05 adaq Target Screens snapshot
241622 06/08/08 19:56 Guy Ron Switching to delta 0 setting
241619 06/08/08 19:08 brads Run 2045 is production
241618 06/08/08 19:06 Guy Ron Back to normal running
241613 06/08/08 18:55 Guy Ron BB Magnet turned off
241611 06/08/08 18:44 brads BB Cerenkov HV Settings
241610 06/08/08 18:43 brads Please do not turn off the BB Cerenkov HVs...
241607 06/08/08 17:35 bogdanw BBite magnet
241606 06/08/08 17:34 adaq Tools and HV status
241601 06/08/08 17:18 Lindgren Shift Summary
241600 06/08/08 15:31 Lindgren Beam Down
241596 06/08/08 13:52 brads Restarted auto-snap Target GUI too
241593 06/08/08 13:36 brads Restarted Tools on auto-snapshot machine
241592 06/08/08 13:20 riordan Automatic Hall A Tools Snapshot Problem
241589 06/08/08 12:23 adaq Radiation and Ozone snapshot
241588 06/08/08 12:23 adaq Hall A Tools snapshot
241587 06/08/08 12:23 adaq BigBox Magnet snapshot
241586 06/08/08 12:23 adaq Beam Position and Current snapshot
241585 06/08/08 12:23 adaq Beam Energy in A and C snapshot
241584 06/08/08 12:20 Lindgren Updated Spot
241583 06/08/08 12:17 Lindgren Hall A Tools and BPM/BCM
241580 06/08/08 12:09 Lindgren Midshift Run Summary
241579 06/08/08 12:05 adaq Target Screens snapshot
241578 06/08/08 11:35 Lindgren Longer runs
241571 06/08/08 08:49 yez Target ESR no value read out, I restarted tgtGUI and it worked.
241570 06/08/08 08:46 Lindgren Day shift starting/Target
241567 06/08/08 08:05 xqian owl shift summary
241565 06/08/08 08:05 yez Target Screen(2)
241564 06/08/08 08:04 yez Target Screen at the Begining of Day Shift
241561 06/08/08 07:34 Moffit Forgot to change LHRS angle in EPICS.
241558 06/08/08 06:38 Xin K1 delta=-2 HV screen shot
241557 06/08/08 06:35 Xin RUN 2031
241554 06/08/08 06:31 adaq RUN 2030 all blocks on K1 delta=-2
241551 06/08/08 05:52 Moffit Moved LHRS to 57.5 deg
241546 06/08/08 04:38 Moffit Target to 6cm Dummy
241544 06/08/08 04:22 adaq Radiation and Ozone snapshot
241543 06/08/08 04:22 adaq Hall A Tools snapshot
241542 06/08/08 04:22 adaq BigBox Magnet snapshot
241541 06/08/08 04:22 adaq Beam Position and Current snapshot
241540 06/08/08 04:22 adaq Beam Energy in A and C snapshot
241539 06/08/08 04:05 adaq Target Screens snapshot
241531 06/08/08 01:54 Moffit Run 2023 using test3 config - no roc9
241527 06/08/08 01:18 brads Turned on all Cerenkov HVs
Saturday
241521 06/07/08 23:53 A.Rakhman Target Screens
241520 06/07/08 23:52 khem shift summary
241519 06/07/08 23:44 khem bpm snap shot
241512 06/07/08 21:39 A.Rakhman Target Screen
241511 06/07/08 21:39 khem snap shot halla tool/ current
241509 06/07/08 21:23 Khem MCC called, should be beam soon.
241508 06/07/08 21:14 brads Controlled Access activity
241507 06/07/08 20:22 adaq Radiation and Ozone snapshot
241506 06/07/08 20:22 adaq Hall A Tools snapshot
241505 06/07/08 20:22 adaq BigBox Magnet snapshot
241504 06/07/08 20:22 adaq Beam Position and Current snapshot
241503 06/07/08 20:22 adaq Beam Energy in A and C snapshot
241501 06/07/08 20:05 adaq Target Screens snapshot
241500 06/07/08 17:55 brads Tech/ Left Q1 tripped, OK now
241498 06/07/08 16:05 bogdanw shift summary
241496 06/07/08 15:55 brads Wood shielding in BB summary
241495 06/07/08 15:38 bogdanw current 1 mkA for DCh. test
241490 06/07/08 15:25 bogdanw Beam 0.9 mkA for DCh test
241483 06/07/08 14:14 xiaohui hv map
241482 06/07/08 14:13 xiaohui Acceptance check
241481 06/07/08 14:12 xiaohui BB shower hit map
241479 06/07/08 13:43 Yi Qiang Target in 6 cm LH2
241477 06/07/08 13:37 bogdanw kin#2 delta = +2%
241474 06/07/08 13:23 Yi Qiang Spot++ of run 2008
241470 06/07/08 13:12 Yi Qiang Target to Carbon
241467 06/07/08 12:38 Yi Qiang Target to Al Dummy
241465 06/07/08 12:22 adaq Radiation and Ozone snapshot
241464 06/07/08 12:22 adaq Hall A Tools snapshot
241463 06/07/08 12:22 adaq BigBox Magnet snapshot
241462 06/07/08 12:22 adaq Beam Position and Current snapshot
241461 06/07/08 12:22 adaq Beam Energy in A and C snapshot
241460 06/07/08 12:05 adaq Target Screens snapshot
241457 06/07/08 11:40 bogdanw beam position and energy
241456 06/07/08 11:39 bogdanw ozone and cryo
241455 06/07/08 11:39 bogdanw HV HRS-L and BB
241454 06/07/08 11:38 bogdanw Tools and beam
241451 06/07/08 10:47 ellie Mistake on End of Run 2002/2003
241448 06/07/08 10:11 bogdanw raster for run 2002
241443 06/07/08 08:23 Yi Qiang Target screen at start of the shift
241442 06/07/08 08:10 R. Michaels partial recovery of fan
241438 06/07/08 08:00 V. Sulkosky Owl Shift Summary
241435 06/07/08 07:24 Moffit Some source EPICS variables wide open.
241430 06/07/08 05:51 V. Sulkosky LHRS HV
241425 06/07/08 04:41 R. Michaels checklist and FPP HV
241424 06/07/08 04:22 adaq Radiation and Ozone snapshot
241423 06/07/08 04:21 adaq Hall A Tools snapshot
241422 06/07/08 04:21 adaq BigBox Magnet snapshot
241421 06/07/08 04:21 adaq Beam Position and Current snapshot
241420 06/07/08 04:21 adaq Beam Energy in A and C snapshot
241419 06/07/08 04:05 adaq Target Screens snapshot
241416 06/07/08 03:47 V. Sulkosky BB HV Map K2 D0%
241415 06/07/08 03:40 V. Sulkosky Acceptance Check K2 D0%
241414 06/07/08 03:37 V. Sulkosky Shower Hit Map 2 D0%
241413 06/07/08 03:31 V. Sulkosky Hall A Tools K2 D0%
241409 06/07/08 03:06 R. Michaels target to 6cm Hydrogen
241407 06/07/08 03:00 V. Sulkosky Spot 1991
241405 06/07/08 02:46 R. Michaels target to 6 cm Al dummy
241401 06/07/08 01:53 R. Michaels target fan is dead
241396 06/07/08 00:18 V. Sulkosky Spot++ 1987
241395 06/07/08 00:11 V. Sulkosky Hall A Tools K2 -2%
Friday
241391 06/06/08 23:59 E.Chudakov Swing shift summary
241390 06/06/08 23:55 A.Rakhman Target Screen at the end of Shift - 2
241389 06/06/08 23:54 A.Rakhman Target Screens at the end of Shift- 1
241386 06/06/08 23:00 xiaohui Fpp asymmetry (run 1984)
241383 06/06/08 22:40 xiaohui Spot++ 1981
241382 06/06/08 22:29 xiaohui Acceptance check
241381 06/06/08 22:19 Nikos Kin. 2 (d=-2) Al run
241380 06/06/08 22:14 xiaohui Start K2 D-2% production from 1984
241373 06/06/08 21:52 E.Chudakov Beam back at about 21:35
241372 06/06/08 21:52 Guy Ron BB HV Map K2 D-2
241371 06/06/08 21:51 Guy Ron BB Hit Map K2 D=-2
241370 06/06/08 21:44 A.Rakhman Target Screen Shots - 2
241369 06/06/08 21:42 A. Rakhman Target Screen Shots - 1
241367 06/06/08 20:50 Nikos beam status
241366 06/06/08 20:33 Nikos BNA since Tuesday night
241365 06/06/08 20:21 adaq Radiation and Ozone snapshot
241364 06/06/08 20:21 adaq Hall A Tools snapshot
241363 06/06/08 20:21 adaq BigBox Magnet snapshot
241362 06/06/08 20:21 adaq Beam Position and Current snapshot
241361 06/06/08 20:21 adaq Beam Energy in A and C snapshot
241360 06/06/08 20:07 brads Correction to HALOG #241353
241359 06/06/08 20:05 adaq Target Screens snapshot
241355 06/06/08 19:43 Brad/Bob Dropped ROC10, DAQ back up.
241353 06/06/08 19:23 brads DAQ problem update
241335 06/06/08 18:13 E.Chudakov Various problems
241319 06/06/08 17:39 Higinbotham Pockel Cell Problem, Return to Q2=0.3
241310 06/06/08 16:53 E.Chudakov Kinematics 2, -2%
241304 06/06/08 16:05 bogdanw shift summary
241300 06/06/08 15:33 Holt 6-cm LH2 target in beam
241299 06/06/08 15:33 brads BB right scint. HVs turned on during Access
241297 06/06/08 15:28 R. Michaels Qasy for recent runs
241296 06/06/08 15:18 R. Michaels bcm u10 broken ?
241294 06/06/08 14:51 bogdanw BB is for kin#1 delta +2%
241293 06/06/08 14:46 bogdanw Tools for kin#1 delta +2%
241292 06/06/08 14:44 Holt 6cm dummy target in beam
241291 06/06/08 14:17 Higinbotham Wood Added
241290 06/06/08 14:16 jpchen target fan and HP heater.
241285 06/06/08 13:55 Higinbotham Adding Wood To BigBite
241282 06/06/08 12:50 Holt more target fan trouble
241280 06/06/08 12:36 bogdanw BB with 0.9 mkA
241278 06/06/08 12:25 adaq BB wire chamber, 1.2 mkA
241276 06/06/08 12:21 adaq Radiation and Ozone snapshot
241275 06/06/08 12:21 adaq Hall A Tools snapshot
241274 06/06/08 12:21 adaq BigBox Magnet snapshot
241273 06/06/08 12:21 adaq Beam Position and Current snapshot
241272 06/06/08 12:21 adaq Beam Energy in A and C snapshot
241271 06/06/08 12:20 Holt 6 cm LH2 target in beam
241267 06/06/08 12:05 adaq Target Screens snapshot
241265 06/06/08 11:59 Holt carbon target in
241261 06/06/08 11:10 fansler Right arm N2 Circuits
241256 06/06/08 09:47 Holt target fan speed
241255 06/06/08 09:46 Holt target screens
241254 06/06/08 09:44 bogdanw spot ++ for run 1940
241253 06/06/08 09:42 bogdanw beam position and energy
241252 06/06/08 09:41 bogdanw BBite magnet
241251 06/06/08 09:40 bogdanw Tools and beam
241246 06/06/08 08:02 V. Sulkosky Owl Shift Summary
241242 06/06/08 07:35 Moffit Target screens
241239 06/06/08 07:06 V. Sulkosky LHRS HV
241236 06/06/08 06:20 Moffit fan frequency fluctuations
241233 06/06/08 05:42 Moffit target: large fluctuations in fan frequency
241228 06/06/08 04:21 adaq Radiation and Ozone snapshot
241227 06/06/08 04:21 adaq Hall A Tools snapshot
241226 06/06/08 04:21 adaq BigBox Magnet snapshot
241225 06/06/08 04:20 adaq Beam Position and Current snapshot
241224 06/06/08 04:20 adaq Beam Energy in A and C snapshot
241221 06/06/08 04:04 adaq Target Screens snapshot
241218 06/06/08 03:32 V. Sulkosky Spot++ 1930
241215 06/06/08 02:43 V. Sulkosky BB HV Map K1 D0%
241214 06/06/08 02:26 V. Sulkosky Acceptance Check
241213 06/06/08 02:18 V. Sulkosky BB Shower Blocks K1 D0%
241212 06/06/08 02:14 V. Sulkosky Hall A Tools K1 0D
241208 06/06/08 01:49 Moffit Moved to Loop3 6cm LH2
241205 06/06/08 01:22 Moffit Target to Al Dummy position
241204 06/06/08 01:21 Moffit Fan frequency transient while moving to Al Dummy
241200 06/06/08 00:26 V. Sulkosky BB HV Trip
241196 06/06/08 00:21 jpchen Swing shift summary
241195 06/06/08 00:11 jpchen Fan speed rasied tp 25% (60Hz).
Thursday
241194 06/05/08 23:56 jpchen BB HVs
241193 06/05/08 23:56 jpchen BB HVs
241192 06/05/08 23:47 jpchen Spot++
241191 06/05/08 23:43 yez Target Screen (2)
241190 06/05/08 23:42 yez Target Screen at the end of Swing shift
241187 06/05/08 23:16 yez Loop3 Target Presure value Should read from...
241178 06/05/08 20:20 adaq Radiation and Ozone snapshot
241177 06/05/08 20:20 adaq Hall A Tools snapshot
241176 06/05/08 20:20 adaq BigBox Magnet snapshot
241175 06/05/08 20:20 adaq Beam Position and Current snapshot
241174 06/05/08 20:20 adaq Beam Energy in A and C snapshot
241171 06/05/08 20:04 adaq Target Screens snapshot
241160 06/05/08 16:08 yez Target Screen(2)
241159 06/05/08 16:07 yez Target Screen at the begining of Swing Shift
241158 06/05/08 16:04 bogdanw shift summary
241153 06/05/08 14:17 xiaohui HV map
241151 06/05/08 14:11 bogdanw Tools and BB for kin#1 -2%
241150 06/05/08 14:10 xiaohui Acceptance check
241149 06/05/08 14:08 xiaohui BB shower blocks for K1 D-2%
241143 06/05/08 13:37 Holt 6-cm LH2 target in beam
241140 06/05/08 13:18 brads HRS-L at 57.5 degree
241137 06/05/08 12:39 Holt 6-cm dummy
241135 06/05/08 12:23 jpchen loop 3 fan
241134 06/05/08 12:20 adaq Radiation and Ozone snapshot
241133 06/05/08 12:20 adaq Hall A Tools snapshot
241132 06/05/08 12:20 adaq BigBox Magnet snapshot
241131 06/05/08 12:20 adaq Beam Position and Current snapshot
241130 06/05/08 12:20 adaq Beam Energy in A and C snapshot
241129 06/05/08 12:04 adaq Target Screens snapshot
241128 06/05/08 11:54 Holt erratic fan speed
241125 06/05/08 11:32 Holt target fan speed alarm
241122 06/05/08 10:58 Holt target screens
241121 06/05/08 10:56 bogdanw spot++ for 1904
241120 06/05/08 10:54 bogdanw crygen and ozone plots
241119 06/05/08 10:53 bogdanw BPM and energy plots
241118 06/05/08 10:52 bogdanw Tools and beam
241117 06/05/08 10:51 brads Possible raster glitch -- OK now but watch for problems
241111 06/05/08 08:01 V. Sulkosky Owl Shift Summary
241106 06/05/08 06:48 V. Sulkosky LHRS VDC HV Tripped
241103 06/05/08 06:19 V. Sulkosky Beam Off
241100 06/05/08 06:13 V. Sulkosky Spot++ 1898
241097 06/05/08 04:40 V. Sulkosky Beam Back
241095 06/05/08 04:34 V. Sulkosky No Beam
241093 06/05/08 04:20 adaq Radiation and Ozone snapshot
241092 06/05/08 04:20 adaq Hall A Tools snapshot
241091 06/05/08 04:19 adaq BigBox Magnet snapshot
241090 06/05/08 04:19 adaq Beam Position and Current snapshot
241089 06/05/08 04:19 adaq Beam Energy in A and C snapshot
241088 06/05/08 04:10 B. Hahn Target screens
241085 06/05/08 04:04 adaq Target Screens snapshot
241078 06/05/08 01:31 Moffit 1891 is junk
241073 06/05/08 01:22 V. Sulkosky Shower Blocks K2 +2D
241072 06/05/08 01:19 V. Sulkosky Spot++ 1890
241071 06/05/08 01:17 V. Sulkosky Run 1890 Starts K2 +2D
241070 06/05/08 01:15 V. Sulkosky HV Map K2 +2D
241069 06/05/08 01:11 V. Sulkosky Acceptance Check
241068 06/05/08 01:00 V. Sulkosky Hall A Tools +2%
241063 06/05/08 00:02 Moffit Run 1887 had raster off.
Wednesday
241060 06/04/08 23:58 Ole Hansen Swing shift summary
241057 06/04/08 23:44 M.Shabestari Target ----> Carbon
241056 06/04/08 23:42 Ole Hansen End delta = 0, now pointing test
241045 06/04/08 21:05 M.Shabestari Target ----> LH2
241043 06/04/08 20:39 M.Shabestari Target ---> Al Dummy
241041 06/04/08 20:26 Ole Hansen Beam coming back. Dummy target in
241040 06/04/08 20:19 adaq Radiation and Ozone snapshot
241039 06/04/08 20:19 adaq Hall A Tools snapshot
241038 06/04/08 20:19 adaq BigBox Magnet snapshot
241037 06/04/08 20:19 adaq Beam Position and Current snapshot
241036 06/04/08 20:19 adaq Beam Energy in A and C snapshot
241035 06/04/08 20:04 adaq Target Screens snapshot
241034 06/04/08 16:13 Ole Hansen No beam
241031 06/04/08 15:23 bogdanw mid shift summary
241030 06/04/08 12:19 adaq Radiation and Ozone snapshot
241029 06/04/08 12:19 adaq Hall A Tools snapshot
241028 06/04/08 12:19 adaq BigBox Magnet snapshot
241027 06/04/08 12:19 adaq Beam Position and Current snapshot
241026 06/04/08 12:19 adaq Beam Energy in A and C snapshot
241025 06/04/08 12:04 adaq Target Screens snapshot
241024 06/04/08 12:04 Holt target charts
241023 06/04/08 12:03 xiaohui BB.ts.sh.e vs BB.ts.ps.e
241022 06/04/08 12:03 Holt target screens-1
241012 06/04/08 09:43 adaq spot++
241011 06/04/08 09:24 adaq Ozone and crygen
241010 06/04/08 09:13 adaq Beam parameters
241009 06/04/08 09:12 adaq BB magnet
241008 06/04/08 09:11 adaq Tools and HV at shift start
241002 06/04/08 08:07 V. Sulkosky Owl Shift Summary
241001 06/04/08 07:48 V. Sulkosky Spot++ 1874
241000 06/04/08 07:46 V. Sulkosky BigBite Power Supply GUI
240996 06/04/08 06:46 V. Sulkosky No Beam
240992 06/04/08 05:46 V. Sulkosky Beam back
240990 06/04/08 05:30 V. Sulkosky No Beam
240986 06/04/08 04:47 EIP FPP gas flow
240985 06/04/08 04:39 B. Hahn Target Screens
240984 06/04/08 04:19 adaq Radiation and Ozone snapshot
240983 06/04/08 04:19 adaq Hall A Tools snapshot
240982 06/04/08 04:19 adaq BigBox Magnet snapshot
240981 06/04/08 04:19 adaq Beam Position and Current snapshot
240980 06/04/08 04:18 adaq Beam Energy in A and C snapshot
240979 06/04/08 04:10 V. Sulkosky BB Shower HV 17.7 Turned Off
240977 06/04/08 04:04 adaq Target Screens snapshot
240975 06/04/08 03:40 V. Sulkosky HV Map K2 D0
240974 06/04/08 03:37 xiaohui Acceptance check & start production from 1868
240973 06/04/08 03:35 xiaohui BB shower blocks K2 D0
240972 06/04/08 03:32 V. Sulkosky Hall A Tools Kinematic 2 (0%)
240964 06/04/08 02:46 B. Hahn Target to 6 cm LH2
240960 06/04/08 02:17 B. Hahn Target to Dummy
240954 06/04/08 00:38 V. Sulkosky Raster Check for 1860
240951 06/04/08 00:00 Boeglin Shift Summary
Tuesday
240945 06/03/08 22:54 Boeglin BB magnet screen shot
240944 06/03/08 22:54 Boeglin BB HV screen shot
240943 06/03/08 22:53 Boeglin BPM screen shot
240942 06/03/08 22:52 Boeglin Spot ++ screen shot
240941 06/03/08 22:51 Armstrong Target screens
240940 06/03/08 22:51 Boeglin Hall A Tools
240933 06/03/08 20:18 adaq Radiation and Ozone snapshot
240932 06/03/08 20:18 adaq Hall A Tools snapshot
240931 06/03/08 20:18 adaq BigBox Magnet snapshot
240930 06/03/08 20:18 adaq Beam Position and Current snapshot
240929 06/03/08 20:18 adaq Beam Energy in A and C snapshot
240926 06/03/08 20:04 adaq Target Screens snapshot
240917 06/03/08 17:37 Armstrong Target screens
240914 06/03/08 16:52 xiaohui HV map
240913 06/03/08 16:46 xiaohui Acceptance check
240912 06/03/08 16:45 xiaohui BB shower blocks K2 D-2
240908 06/03/08 16:13 brads BigBite angle -> 30.4 degrees
240907 06/03/08 16:03 bogdanw shift summary
240901 06/03/08 15:30 Moffit EDTM to 12.5kHz during 1844
240900 06/03/08 12:18 adaq Radiation and Ozone snapshot
240899 06/03/08 12:18 adaq Hall A Tools snapshot
240898 06/03/08 12:18 adaq BigBox Magnet snapshot
240897 06/03/08 12:18 adaq Beam Position and Current snapshot
240896 06/03/08 12:18 adaq Beam Energy in A and C snapshot
240895 06/03/08 12:04 adaq Target Screens snapshot
240894 06/03/08 10:34 bogdanw Tools and BB HV
240893 06/03/08 10:20 adaq HRS-L angle
240892 06/03/08 10:05 Holt Target screens
240891 06/03/08 09:30 bogdanw FPP door positions
240889 06/03/08 08:39 bogdanw BPMs and cryolevels
240888 06/03/08 08:38 bogdanw HV screens
240887 06/03/08 08:37 bogdanw Tools and beam
240886 06/03/08 08:35 bogdanw ozone monitor beam energy
240882 06/03/08 08:07 A. Camsonne Owl shift summary
240877 06/03/08 06:52 A. Camsonne BB power supply
240876 06/03/08 06:50 A. Camsonne BPM/BCM
240875 06/03/08 06:49 A. Camsonne BB HV
240874 06/03/08 06:49 A. Camsonne Hall A tools
240873 06/03/08 06:44 A. Camsonne Raster check run 1840
240868 06/03/08 04:21 P. Decowski Target charts
240867 06/03/08 04:20 P. Decowski Target screens
240864 06/03/08 04:18 adaq Radiation and Ozone snapshot
240863 06/03/08 04:18 adaq Hall A Tools snapshot
240862 06/03/08 04:18 adaq BigBox Magnet snapshot
240861 06/03/08 04:18 adaq Beam Position and Current snapshot
240860 06/03/08 04:18 adaq Beam Energy in A and C snapshot
240859 06/03/08 04:04 adaq Target Screens snapshot
240854 06/03/08 01:56 adaq FPP adc shift not BB
240853 06/03/08 01:36 eip BB adc shift ?
240850 06/03/08 01:07 A. Camsonne Beam back
240849 06/03/08 00:48 A. Camsonne Beam off machine magnet trip
240848 06/03/08 00:03 M.Shabestari Swing Shift Summary
Monday
240844 06/02/08 23:45 M.Shabestari BB HV shower map
240843 06/02/08 23:17 M.Shabestari No beam for a while
240838 06/02/08 21:35 M.Shabestari BB HV tripped due to beam trip
240835 06/02/08 21:31 Chiranjib TARGET ALARM
240832 06/02/08 20:37 M.Shabestari Ozone Monitor Snapshot
240831 06/02/08 20:29 Nikos Progress of Kin. # 8 (d=+2%)
240830 06/02/08 20:17 adaq Radiation and Ozone snapshot
240829 06/02/08 20:17 adaq Hall A Tools snapshot
240828 06/02/08 20:17 adaq BigBox Magnet snapshot
240827 06/02/08 20:17 adaq Beam Position and Current snapshot
240826 06/02/08 20:17 adaq Beam Energy in A and C snapshot
240825 06/02/08 20:03 adaq Target Screens snapshot
240822 06/02/08 19:26 M.Shabestari BigBox Screens
240821 06/02/08 19:19 Chiranjib TARGET SCREENS
240818 06/02/08 18:47 M.Shabestari Hall A Tools snapshot
240817 06/02/08 18:44 M.Shabestari Spot++ for run 1827
240816 06/02/08 18:29 M.Shabestari Beam Energy snapshot
240815 06/02/08 18:28 M.Shabestari BB HV tripped.
240810 06/02/08 17:24 M.Shabestari Beam is back, data-taking resumed.
240809 06/02/08 17:21 M.Shabestari BPM and BCM snapshots
240807 06/02/08 17:03 adaq No beam - MCC's software problem
240802 06/02/08 16:08 jpchen Day shift summary
240801 06/02/08 16:00 Armstrong FastFeedback off
240800 06/02/08 15:50 brads Controlled Access activity
240799 06/02/08 15:39 Armstrong beam back, data taking resumed
240797 06/02/08 15:33 Armstrong no beam - IOC problem
240792 06/02/08 14:26 jpchen ROC5 and ROC 8 problem
240787 06/02/08 12:26 Armstrong Target fan speed fluctuations
240786 06/02/08 12:17 adaq Radiation and Ozone snapshot
240785 06/02/08 12:17 adaq Hall A Tools snapshot
240784 06/02/08 12:17 adaq BigBox Magnet snapshot
240783 06/02/08 12:17 adaq Beam Position and Current snapshot
240782 06/02/08 12:17 adaq Beam Energy in A and C snapshot
240781 06/02/08 12:03 adaq Target Screens snapshot
240778 06/02/08 11:38 Armstrong Target screens
240777 06/02/08 11:30 A. Camsonne Detector testing for CSR thickness measurement
240775 06/02/08 10:34 Armstrong back to 6 cm lH2 target
240774 06/02/08 10:32 R. Michaels auto_abu repaired (BNA problem)
240772 06/02/08 10:18 jpchen BNA
240770 06/02/08 09:52 Armstrong move to Al dummy
240768 06/02/08 08:55 xiaohui Two more shower blocks on
240763 06/02/08 08:09 A. Camsonne Owl shift summary
240761 06/02/08 07:52 A. Camsonne Spot for run 1810
240760 06/02/08 07:48 A. Camsonne BPM/BCM chart
240759 06/02/08 07:33 Guy Ron BB HV Map K8 +2%
240757 06/02/08 07:32 Guy Ron K8 +2% Acctepance Check
240750 06/02/08 06:55 A. Camsonne Raster size changed
240749 06/02/08 06:54 P. Decowski Hydrogen 6cm target
240748 06/02/08 06:46 P. Decowski Al target in
240747 06/02/08 06:46 A. Camsonne Changed left arm momentum to 0.8948 GeV/c
240745 06/02/08 06:37 A. Camsonne Spot for run 1809
240742 06/02/08 06:29 A. Camsonne Beam position for pointing run 1808
240740 06/02/08 06:22 P. Decowski Carbon target in
240739 06/02/08 06:18 A. Camsonne Kin 8 at 0 % complete
240737 06/02/08 06:11 A. Camsonne Spot run 1807
240736 06/02/08 06:09 P. Decowski Target screens
240735 06/02/08 06:08 P. Decowski Target charts
240734 06/02/08 05:48 A. Camsonne Hall A tools Kin 8 at 0%
240729 06/02/08 04:17 adaq Radiation and Ozone snapshot
240728 06/02/08 04:17 adaq Hall A Tools snapshot
240727 06/02/08 04:17 adaq BigBox Magnet snapshot
240726 06/02/08 04:17 adaq Beam Position and Current snapshot
240725 06/02/08 04:17 adaq Beam Energy in A and C snapshot
240724 06/02/08 04:03 adaq Target Screens snapshot
240717 06/02/08 01:44 A. Camsonne Beam back
240716 06/02/08 01:39 A. Camsonne No beam
240715 06/02/08 01:13 A. Camsonne ROC 10 reset
240711 06/02/08 00:12 A. Camsonne Beam back
240710 06/02/08 00:06 A. Camsonne No beam for a few minutes
Sunday
240708 06/01/08 23:57 Xin Swing shift summary
240704 06/01/08 22:35 Kalyan Target to H2
240701 06/01/08 22:14 Kalyan Target to Al
240695 06/01/08 20:17 adaq Radiation and Ozone snapshot
240694 06/01/08 20:16 adaq Hall A Tools snapshot
240693 06/01/08 20:16 adaq BigBox Magnet snapshot
240692 06/01/08 20:16 adaq Beam Position and Current snapshot
240691 06/01/08 20:16 adaq Beam Energy in A and C snapshot
240690 06/01/08 20:03 adaq Target Screens snapshot
240689 06/01/08 19:50 A. Rakhman run 1794 ROC 8 crash
240687 06/01/08 19:30 Kalyan BB HV trip
240684 06/01/08 19:01 Kalyan Target screen shots
240683 06/01/08 18:53 xin right arm ioc reboot
240680 06/01/08 18:28 xiaohui FPP efficiency is recovering
240679 06/01/08 18:12 xin BB HV screen shot
240678 06/01/08 18:11 xiaohui Acceptance check
240677 06/01/08 18:09 xiaohui BB shower blocks K8 D0
240674 06/01/08 17:44 A. Rakhman run 1790
240669 06/01/08 16:48 Nikos Kin.#8 update
240666 06/01/08 16:03 M.Shabestari Target Screens
240665 06/01/08 16:03 M.Shabestari Day Shift Summary
240661 06/01/08 14:46 M.Shabestari Change in Front & Back chambers hits
240658 06/01/08 13:27 folts tech/ detector gas alarm
240657 06/01/08 13:26 M.Shabestari Low pressure problem fixed
240656 06/01/08 13:21 M.Shabestari Low pressure alarm
240655 06/01/08 13:14 M.Shabestari Spot++ for run 1784 (LH2 6cm)
240650 06/01/08 12:18 M.Shabestari BCM &BPM Stripcharts
240649 06/01/08 12:16 adaq Radiation and Ozone snapshot
240648 06/01/08 12:16 adaq Hall A Tools snapshot
240647 06/01/08 12:16 adaq BigBox Magnet snapshot
240646 06/01/08 12:16 adaq Beam Position and Current snapshot
240645 06/01/08 12:16 adaq Beam Energy in A and C snapshot
240644 06/01/08 12:10 M.Shabestari BigBox Screens
240643 06/01/08 12:03 adaq Target Screens snapshot
240642 06/01/08 11:42 M.Shabestari Target Screens
240639 06/01/08 10:42 Nikos progress of Kin. #8
240638 06/01/08 10:38 Shabestari Hall A Tools Snapshot
240635 06/01/08 08:59 M.Shabestari BPM and BCM Snapshots
240632 06/01/08 08:01 A. Camsonne Owl shift summary
240630 06/01/08 07:48 A. Camsonne Beam is back
240628 06/01/08 07:41 A. Camsonne BB power supply
240626 06/01/08 07:35 A. Camsonne MCC will do a spot move at 7:35
240625 06/01/08 07:24 A. Camsonne BB HV
240624 06/01/08 07:23 A. Camsonne BPM / BCM
240623 06/01/08 07:21 A. Camsonne Hall A tools Kin 8 -2%
240620 06/01/08 06:39 Piotr Decows Target charts
240619 06/01/08 06:35 Piotr Decows `Target screens
240614 06/01/08 04:16 adaq Radiation and Ozone snapshot
240613 06/01/08 04:16 adaq Hall A Tools snapshot
240612 06/01/08 04:16 adaq BigBox Magnet snapshot
240611 06/01/08 04:16 adaq Beam Position and Current snapshot
240610 06/01/08 04:16 adaq Beam Energy in A and C snapshot
240609 06/01/08 04:03 adaq Target Screens snapshot
240606 06/01/08 03:14 A. Camsonne BTA oddity
240605 06/01/08 02:54 A. Camsonne Beam fine after resuming
240602 06/01/08 02:32 A. Camsonne Beam down for a few minutes
240597 06/01/08 00:21 Nikos progress of Kin.#8 (d=-2%)
240596 06/01/08 00:19 A. Camsonne Raster check run 1773
240595 06/01/08 00:05 M.Shabestari Swing Shift Summary